Circular No. 8336

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004bq IN ESO 597-G32 Further to IAUC 8335, M. Moore and W. Li report the LOSS discovery,

on KAIT images taken on May 5.48 and 6.45 UT, of an apparent supernova (mag ∼ 17.3) located at α = 20^{h}43^{m}58^{s}.28, δ = −20^{o}40′25′′.7 (equinox 2000.0), which is 2′′.9 west and 8′′.3 north of the nucleus of ESO 597-G32. A KAIT image taken on 2003 Oct. 28.22 UT showed nothing at this position (limiting mag ∼ 19.0).

COMETS C/2004 B2, C2, H4, AND H5 (SOHO) Further to IAUC 8334, K. Battams reports measurements for ad-

ditional Meyer-group comets found on SOHO C2 images by R. Kracht (C/2004 B2 and H4), X. Leprette (C/2004 C2), and himself (C/2004 H5). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2004-J21.

Comet  2004 UT  α_{2000}  δ_{2000}
C/2004 B2 Jan. 28.913  20^{h}47.^{m}9  −17^{o}09′
C/2004 C2 Feb. 4.963  21 16.7  −15 10
C/2004 H4 Apr. 16.738  1 45.4  +11 59
C/2004 H5  30.588  2 37.4  +16 13
2003 YT_{1}

M. C. Nolan, E. S. Howell, and A. A. Hine, National Astronomy and Ionosphere Center, report: “Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm), obtained on May 1–4, show that 2003 YT_{1} (cf. MPEC 2003-Y30) is a binary system. Preliminary estimates of average diameters, based on range extents at 15-m and 30-m resolution, are 1000 and 180 m. The primary has an irregular shape and a rotation period no longer than 2.6 hr. The orbit appears to be eccentric, with a period near 30 hr. The secondary has a rotation period no longer than 6 hr, inconsistent with synchronous rotation.”

A. Galád, Š. Gajdoš, L. Kornoš and J. Vilagi, Modra Observatory; and P. Pravec, Ondřejov Observatory, report that photometric observations — obtained on three nights during Apr. 21–29 at Modra — show a primary rotation period of 2.343 ± 0.001 hr and a lightcurve amplitude of 0.16 mag, which is indicative of a roughly spheroidal shape. A mutual attenuation event 0.15 mag deep and 0.7 hr long was detected around Apr. 22.99 UT.

2004 May 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green